OpenStudy (anonymous):

What is the value of a? -17/a = -17

OpenStudy (♪chibiterasu):

We just have to isolate a. We need to move the a to the right. Do you know how to do that?

OpenStudy (anonymous):

kinda

OpenStudy (♪chibiterasu):

Ok, so how do you think we have to move the 'a' from -17/a to the right with -17?

OpenStudy (anonymous):

im not a big fan of fractions

OpenStudy (anonymous):

-17/-17 ?

OpenStudy (♪chibiterasu):

And that's why we get rid of them, To get rid of it, we just have to multply a on both sides. -17/a is -17 divided by a. When we multiply a into it, it cancels out and leaves us with -17. Now we multiply a on the other side and it gives us -17a.

OpenStudy (♪chibiterasu):

-17/a * a = -17 * a -17 = -17a

OpenStudy (♪chibiterasu):

Does that make sense?

OpenStudy (anonymous):

Yea it does

OpenStudy (♪chibiterasu):

Now we have to keep isolating a. How do you think we can move the -17 in -17a to the other side of the equation?

OpenStudy (anonymous):

I just lost my answer on what i was going to type

OpenStudy (♪chibiterasu):

Hmm?

OpenStudy (anonymous):

hmmm

OpenStudy (♪chibiterasu):

The things you type disappear sometimes, the site has a lot of lag.

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

i have to leave in a little bit

OpenStudy (♪chibiterasu):

Ok, well. Remember that we need to cancel out. What is the opposite of multiplication?

OpenStudy (anonymous):

division

OpenStudy (♪chibiterasu):

So what do we do with -17 in the -17a?

OpenStudy (anonymous):

divide it

OpenStudy (♪chibiterasu):

Yes. What do we get when we divide -17 on both sides of the equation?

OpenStudy (anonymous):

.

OpenStudy (anonymous):

you divide it with a right

OpenStudy (♪chibiterasu):

No. We need to isolate a, so it's going to be: -17/-17 = (-17a)/-17

OpenStudy (♪chibiterasu):

what do we get

OpenStudy (♪chibiterasu):

When you're dividing a number by itself what do you get

OpenStudy (anonymous):

-17 or 0

OpenStudy (anonymous):

or 1

OpenStudy (♪chibiterasu):

Yes, 1. Sooooo.....

OpenStudy (♪chibiterasu):

-17/-17 = ? -17a/-17 = ?

OpenStudy (anonymous):

a=-17?

OpenStudy (♪chibiterasu):

No

OpenStudy (♪chibiterasu):

We are dividing both sides by -17

OpenStudy (anonymous):

a=1?

OpenStudy (♪chibiterasu):

Yes!

OpenStudy (anonymous):

gg thx

OpenStudy (anonymous):

i give you a medal thx

OpenStudy (♪chibiterasu):

Thanks xD